# C2-Evo: Co-Evolving Multimodal Data and Model for Self-Improving Reasoning

### v1

@@ 1 Introduction
To address these challenges, we propose a fully automated, adaptive multimodal learning framework (C2-Evo) that jointly evolves both the model and its training data in a closed-loop fashion, with a particular focus on geometric reasoning tasks. Table 1 summarizes the differences between our framework and state-of-the-art methods: Unlike existing methods that rely on static data, our method dynamically adjusts task complexity based on real-time assessments of model performance, ensuring a tighter coupling between model capability and data difficulty throughout the learning process. Specifically, to tackle the challenge (1), we incorporate the process into a cross-modal data evolution loop, where we integrate SKETCHPAD-generated complex geometric diagrams [19] with complex problem synthesis (cf. Figure 1 (a)). GPT-4o is employed to generate auxiliary construction code, then executed via tools (e.g., Jupyter). Formal image descriptions are produced using the Doubao model. Sub-problems (4–10 per image) are then generated following predefined guidelines and composed into complex reasoning questions. These are filtered using GPT-4o to remove misaligned cases, and validated using DeepSeek’s three-step reasoning framework to ensure consistency. This process yields a dataset with semantically aligned visual and textual components. For the challenge (2), we adopt a data-model evolution loop. This loop utilizes data generated from the cross-modal data evolution loop and applies it to iteratively fine-tune the model using SFT and RL. SFT maintains output structure and coherence, while RL improves generalization through rule-based optimization. We introduce a simple error-based filtering method that evaluates sample difficulty via prediction variance over $32$ generations. By selecting samples with an error rate $\geq 0.3$, which are prone to error yet still within the model’s grasp and pose a meaningful challenge, the framework ensures that task difficulty remains aligned with model capability, achieving continuous improvement over iterations (cf. , Figure 1 (c)).
@@ 3.2 Co-Evolving Multimodal Data
Using these principles, we prompt DeepSeek-R1 with the tuples $(F_{t},R_{t},P)$ to generate a diverse set of sub-problems ($q_{1},q_{2},\ldots,q_{m}$, where $m$ typically ranges from $4$ to $10$, depending on the complexity of the image). To mitigate the inherent limitations of formal language in capturing visual content, we incorporate the role-playing strategy from R1-Onevision [51]. By analyzing the differences and commonalities among sub-problems ($q_{1},q_{2},\ldots,q_{m}$), we align them with corresponding sub-image elements to compose challenging geometric reasoning questions. Specifically, by combining the sub-problems ($q_{1},q_{2},\ldots,q_{m}$) with their associated sub-image descriptions $F_{t}$, we prompt DeepSeek-R1 to compose challenging questions. For example, given two sub-problems $q_{1}$ and $q_{2}$ that address different parts of an image, we combine them into a more challenging question $Q_{t+1}$ by leveraging their thematic and contextual connections.
@@ 3.3 Co-Evolving Data and Model
In the $t$-th iteration, we perform SFT followed by RL training using all post-filtered data, resulting in an updated model $\pi_{\theta,RL}^{t+1}$ and a new dataset $\mathcal{D}_{t+1}$ for the next round. To align the difficulty of $\mathcal{D}$ with the current capability of $\pi_{\theta}$, we forward each sample through the model $32$ times and compute the error rate as the proportion of times the model generated a wrong answer. As illustrated in Figure 4, we evaluate the difficulty of the data for each model obtained from RL training. We then retain only the samples with an error rate greater than $0.3$ to form the training set for the next iteration, ensuring that the data complexity remains aligned with model capability.
@@ 4.1 Benchmark
Implementation Details. In our experiments, we adopt two state-of-the-art open-source MLLMs, i.e., Qwen2-VL-2B [45] and Qwen2-VL-7B [45]. For the policy warm-up phase, we employ the LLaMA-Factory [^4] https://github.com/hiyouga/LLaMA-Factory. framework with a batch size of $128$ and a learning rate of $51e-$ 5. For the RL phase, we use the VLM-R1555https://github.com/om-ai-lab/VLM-R1. framework. In the first iteration, we perform $32$ rollouts per question, reducing to $8$ in subsequent iterations. The temperature is set to the default value of $0.9$, and the KL divergence coefficient $\beta$ in Equation 2 is set to 00. All experiments are conducted on $32$ NVIDIA V100-$32$GB GPUs.
We evaluate C2-Evo on several multimodal reasoning benchmarks, including Geo-Sub from Geometry3k-test [32], MathVista [31], and MathVerse [56]. To provide a more comprehensive evaluation of model performance, we construct a specialized test subset by sampling images from Geometry3K-test that require the use of auxiliary lines to solve. This results in a focused benchmark containing $274$ images. When evaluated on the original images, we denote the setting as Geo-Sub. Using the corresponding images with added auxiliary lines, we refer to the setting as Geo-Sub-Aux. Further details can be found in the Supplementary Material.
@@ 4.2 Main Results
C2-Evo-$1^{st}$ 0.6k 45.5 48.7 55.0 57.0 54.0 62.3 33.6
C2-Evo-$2^{nd}$ 0.6k 46.9 50.9 56.0 58.0 56.0 62.7 34.8
C2-Evo-$3^{rd}$ 0.4k 50.9 54.9 59.0 59.0 60.0 63.4 34.9
As shown in Table 2, C2-Evo demonstrates a significant improvement over Qwen2-VL-7B and Qwen2-VL-2B across three benchmarks through three iterations, with geometric reasoning accuracy increasing notably from $40.4$ to $54.9$. The progressive improvement observed across three iterations highlights the effectiveness of our self-improving strategy in enhancing reasoning capabilities. Furthermore, compared to prior reasoning models, our approach achieves superior performance while leveraging less than $1\%$ of the training data, with performance approaching that of GPT-4o.
@@
The Influence of error-rate. As described in Section 3.3, we conduct an extensive parameter analysis over varying values of error-rate, which governs the complexity of data within each iteration. Table 5 shows the performance of models trained solely on data from the current iteration $D_{t}$, compared to those that also incorporate historical data $D_{1,\ldots,t-1}$. Utilizing the complete dataset does not necessarily optimize the model’s reasoning capabilities. Specifically, as the complexity of the training data increases, relying exclusively on error-prone samples results in a decline performance. Based on this analysis, error-rate0.3absent0.3$\geq$ 0.3≥ 0.3 is adopted to as the basis for our final setting.
@@ A.1 Implementation Details
We have introduced auxiliary lines into the Geo-Sub-Aux dataset to facilitate problem solving and improve model alignment during training. Originating from Geometry3k[32], the Geo-Sub dataset comprises $275$ samples. The left image illustrates the structure of the Geo-Sub dataset (cf. , Figure VI left), whereas the right image showcases the Geo-Sub-Aux dataset (cf. , Figure VI right), highlighted by its additional auxiliary lines.
@@
As shown in Figure VII, we also evaluate the difficulty of training data from previous iterations using the model training in the preceding iteration. Our observations are as follows: 1) In the evaluation of the second iteration, the first iteration model $\pi_{\theta,\text{RL}}^{1}$ achieved completely wrong predictions on approximately $8\%$ of the data from the first iteration. 2) Despite the fact that the data generated in the second iteration was more challenging than that of the first iteration (cf. , as evidenced by the longer reasoning length in the Figure $3$, the first iteration model $\pi_{\theta,\text{RL}}^{1}$ was still able to solve about $30\%$ of the problems from the second iteration. 3) In the evaluation of the third iteration, when the second iteration model $\pi_{\theta,\text{RL}}^{2}$ was used to evaluate the data from the first iteration, there is an increase in completely wrong predictions and a decrease in fully correct ones. This suggests that the model begins to exhibit catastrophic forgetting. 4) Even within the second iteration data, around $30\%$ of the questions are completely wrong ,which aligns with the assumption stated in the main text that the data difficulty increases across iterations.

### v2

@@ 1 Introduction
To address these challenges, we propose a fully automated, adaptive multimodal learning framework (C2-Evo) that jointly evolves both the model and its training data in a closed-loop fashion, with a particular focus on geometric reasoning tasks. Table 1 summarizes the differences between our framework and state-of-the-art methods: Unlike existing methods that rely on static data, our method dynamically adjusts task complexity based on real-time assessments of model performance, ensuring a tighter coupling between model capability and data difficulty throughout the learning process. Specifically, to tackle the challenge (1), we incorporate the process into a cross-modal data evolution loop, where we integrate SKETCHPAD-generated complex geometric diagrams [19] with complex problem synthesis (cf. Figure 1 (a)). GPT-4o is employed to generate auxiliary construction code, then executed via tools (e.g., Jupyter). Formal image descriptions are produced using the Doubao model. Sub-problems (4–10 per image) are then generated following predefined guidelines and composed into complex reasoning questions. These are filtered using GPT-4o to remove misaligned cases, and validated using DeepSeek’s three-step reasoning framework to ensure consistency. This process yields a dataset with semantically aligned visual and textual components. For the challenge (2), we adopt a data-model evolution loop. This loop utilizes data generated from the cross-modal data evolution loop and applies it to iteratively fine-tune the model using SFT and RL. SFT maintains output structure and coherence, while RL improves generalization through rule-based optimization. We introduce a simple error-based filtering method that evaluates sample difficulty via prediction variance over 323232 generations. By selecting samples with an error rate 0.3$\geq$ 0.3≥ 0.3, which are prone to error yet still within the model’s grasp and pose a meaningful challenge, the framework ensures that task difficulty remains aligned with model capability, achieving continuous improvement over iterations (cf. , Figure 1 (c)).
@@ 3.2 Co-Evolving Multimodal Data
Using these principles, we prompt DeepSeek-R1 with the tuples $(F_{t},R_{t},P)$ to generate a diverse set of sub-problems ($q_{1},q_{2},\ldots,q_{m}$, where $m$ typically ranges from 444 to 101010, depending on the complexity of the image). To mitigate the inherent limitations of formal language in capturing visual content, we incorporate the role-playing strategy from R1-Onevision [51]. By analyzing the differences and commonalities among sub-problems ($q_{1},q_{2},\ldots,q_{m}$), we align them with corresponding sub-image elements to compose challenging geometric reasoning questions. Specifically, by combining the sub-problems ($q_{1},q_{2},\ldots,q_{m}$) with their associated sub-image descriptions $F_{t}$, we prompt DeepSeek-R1 to compose challenging questions. For example, given two sub-problems $q_{1}$ and $q_{2}$ that address different parts of an image, we combine them into a more challenging question $Q_{t+1}$ by leveraging their thematic and contextual connections.
@@ 3.3 Co-Evolving Data and Model
In the $t$-th iteration, we perform SFT followed by RL training using all post-filtered data, resulting in an updated model $\pi_{\theta,RL}^{t+1}$ and a new dataset $\mathcal{D}_{t+1}$ for the next round. To align the difficulty of $\mathcal{D}$ with the current capability of $\pi_{\theta}$, we forward each sample through the model 323232 times and compute the error rate as the proportion of times the model generated a wrong answer. As illustrated in Figure 4, we evaluate the difficulty of the data for each model obtained from RL training. We then retain only the samples with an error rate greater than 0.30.30.3 to form the training set for the next iteration, ensuring that the data complexity remains aligned with model capability.
@@ 4.1 Benchmark
Implementation Details. In our experiments, we adopt two state-of-the-art open-source MLLMs, i.e., Qwen2-VL-2B [45] and Qwen2-VL-7B [45]. For the policy warm-up phase, we employ the LLaMA-Factory [^4] https://github.com/hiyouga/LLaMA-Factory. framework with a batch size of 128128128 and a learning rate of $1e-5$. For the RL phase, we use the VLM-R1555https://github.com/om-ai-lab/VLM-R1. framework. In the first iteration, we perform 323232 rollouts per question, reducing to 888 in subsequent iterations. The temperature is set to the default value of 0.90.90.9, and the KL divergence coefficient $\beta$ in Equation 2 is set to 0. All experiments are conducted on 323232 NVIDIA V100- [^32] GB GPUs.
We evaluate C2-Evo on several multimodal reasoning benchmarks, including Geo-Sub from Geometry3k-test [32], MathVista [31], and MathVerse [56]. To provide a more comprehensive evaluation of model performance, we construct a specialized test subset by sampling images from Geometry3K-test that require the use of auxiliary lines to solve. This results in a focused benchmark containing 274274274 images. When evaluated on the original images, we denote the setting as Geo-Sub. Using the corresponding images with added auxiliary lines, we refer to the setting as Geo-Sub-Aux. Further details can be found in the Supplementary Material.
@@ 4.2 Main Results
[…]
@@
The Influence of error-rate. As described in Section 3.3, we conduct an extensive parameter analysis over varying values of error-rate, which governs the complexity of data within each iteration. Table 5 shows the performance of models trained solely on data from the current iteration $D_{t}$, compared to those that also incorporate historical data $D_{1,\ldots,t-1}$. Utilizing the complete dataset does not necessarily optimize the model’s reasoning capabilities. Specifically, as the complexity of the training data increases, relying exclusively on error-prone samples results in a decline performance. Based on this analysis, error-rate0.3$\geq$ 0.3≥ 0.3 is adopted to as the basis for our final setting.
@@ A.1 Implementation Details
We have introduced auxiliary lines into the Geo-Sub-Aux dataset to facilitate problem solving and improve model alignment during training. Originating from Geometry3k[32], the Geo-Sub dataset comprises 275275275 samples. The left image illustrates the structure of the Geo-Sub dataset (cf. , Figure VI left), whereas the right image showcases the Geo-Sub-Aux dataset (cf. , Figure VI right), highlighted by its additional auxiliary lines.
@@
[…]